OpenStudy (anonymous):

A manufacturer needs to make a cylinder can that will hold 1.5 liters of liquid. Determine the dimensions of the can that will minimize the amount of material used in itd construction.

OpenStudy (anonymous):

Determine the dimensions of the cylinder given a volume of 1.5 liters. so, 1.5 liters = formula for a cylinder volume 1.5 liters = pi*r^2 (h) minimize the amount of material used so, material used is equivalent to the surface area of the cylinder. area = (2)pi*r^2 + (h) 2pi*r

OpenStudy (anonymous):

we need to convert liters meters^3 1.5L = 0.0015 m^3 0.0015 = pi*r^2 (h) set equal to h h = 0.0015 / (pi*r^2) combine with area equation by replacing h area = (2)pi*r^2 + (h) 2pi*r area = (2)pi*r^2 + 2pi*r (0.0015 / (pi*r^2)) to 'minimize' we need to take the derivative of area with respect to r, set it equal to zero. solve for r. once you have r, plug it into 0.0015 = pi*r^2 (h) and solve for h